# Resolving the Module of Derivations on an $n\times(n+1)$ Determinantal Ring

### v1

@@ 1. Introduction
This paper is concerned with showing that a truncation of a relative bar resolution minimally resolves the module of $k$-linear derivations $\operatorname{Der}_{R\mid k}$ on a determinantal ring $R$ defined by the maximal minors of an $n\times(n+1)$ generic matrix $X$. The minimality of the resolution yields a minimal generating set for $\operatorname{Der}_{R\mid k}$ over $R$, a result which may be found in [bv], produced by a starkly different method.
While minimal generators of modules of derivations on determinantal rings are known, the minimal graded free resolution of $\operatorname{Der}_{R\mid k}$ over $R$ has not been computed. The first order differential operators $D^{1}_{R\mid k}$ on $R$ decompose as $D^{1}_{R\mid k}\cong\operatorname{Der}_{R\mid k}\oplus R$, and so after obtaining the resolution of $\operatorname{Der}_{R\mid k}$, the resolution of $D^{1}_{R\mid k}$ is immediate. Known resolutions of differential operators are few -- see [rc] for the case of low-order differential operators on an isolated hypersurface singularity and [yu] for derivations on a generic hyperplane arrangement, as two examples. Differential operators need not behave the same in positive characteristic as in characteristic zero, so it seems worthwhile to point out that our approach to the minimal graded $R$-free resolution of $\operatorname{Der}_{R\mid k}$ is characteristic-free.
@@
$0\to\operatorname{Der}_{R\mid k}\to R^{n(n+1)}\overset{J^{T}_{R}}{\to}R^{n+1}\to\operatorname{coker}J^{T}_{R}\to 0$
@@
$\ldots\longrightarrow R^{n(n+1)(n+2)}\overset{\small\begin{bmatrix}J^{T}_{R}&\varphi_{2}\\ M_{1,1}&M_{2,0}\end{bmatrix}}{\longrightarrow}R^{2n(n+1)}\overset{\small\begin{bmatrix}\partial_{2}&M_{1,0}\end{bmatrix}}{\longrightarrow}\operatorname{Der}_{R\mid k}\longrightarrow 0,$
@@ 2. Background
$0\longrightarrow Q^{n}\overset{\varphi_{2}}{\longrightarrow}Q^{n+1}\overset{\varphi_{1}}{\longrightarrow}Q\longrightarrow R\longrightarrow 0,$
@@
The ring $R$ is Golod if and only if $\mathcal{A}_{\geq 1}\cdot\mathcal{A}_{\geq 1}\subseteq\mathfrak{m}_{Q}\mathcal{A}$, where $\mathfrak{m}_{Q}$ is the (homogeneous) maximal ideal of $Q$. Similarly, an $S/J$-module $N$ is Golod if and only if its minimal $Q$-free resolution $\mathcal{U}$ is a dg module over $\mathcal{A}$ with structure satisfying that the action of $\mathcal{A}_{\geq 1}$ on $\mathcal{U}$ lands in $\mathfrak{m}_{Q}\mathcal{U}$.
@@
where the bars mean to shift by 1 homological degree.
Define $m_{n}:A^{\otimes n}\to A$ iteratively by $m_{1}=\partial^{A}$ and $m_{n}$ ($n>1$) by the characterization that it satisfies the Stasheff identity
$\sum_{s=1}^{n}\sum_{r+s+t=n}(-1)^{r+st}m_{r+1+t}(1^{\otimes r}\otimes m_{s}\otimes 1^{\otimes t})=0,$
in which $r,t\geq 0$. Define $\mu_{n}:A^{\otimes n-1}\otimes U\to U$ such that analogous identities hold.
@@
$\partial(s\otimes a_{1}\otimes\ldots\otimes a_{p}\otimes u)$ $=\sum_{i=1}^{p}\sum_{j=0}^{p-i}\pm s\otimes a_{1}\otimes\ldots\otimes a_{j}\otimes m_{i}(a_{j+1}\otimes a_{j+i})\otimes a_{j+i+1}\ldots\otimes a_{p}\otimes u$
$+\sum_{i=1}^{p+1}\pm s\otimes a_{1}\otimes\ldots\otimes a_{p-i+1}\otimes\mu_{i}(a_{p-i+2}\otimes\ldots\otimes a_{p}\otimes u).$
Exact signs are suppressed, and they are unimportant to the moral of the story.
@@ A $k$-linear derivation $\delta$ on the polynomial $k$-algebra $k[x_{1},\ldots,x_{s}]$ can be written
where $\alpha_{j}\in k$ and $\dfrac{\partial}{\partial x_{j}}$ is partial differentiation with respect to $x_{j}$. The prototypical example of a derivation that does not look like a derivation is the Euler derivation
$E(f):=\deg(f)\cdot f,$
where $f$ is a homogeneous polynomial. One can express $E\in\operatorname{Der}_{Q\mid k}$ (or in $\operatorname{Der}_{R\mid k}$, where $x_{k,\ell}$ is understood to be an equivalence class) as
$E=\displaystyle\sum_{k=1}^{n}\sum_{\ell=1}^{n+1}x_{k,\ell}\dfrac{\partial}{\partial x_{k,\ell}}.$
@@ 3. Useful Lemmas
(3) $\sum_{r\neq i}(-1)^{r}\dfrac{\partial F_{r}}{\partial x_{j,i}}x_{j,r}=(-1)^{i+1}F_{i},$
@@
To see (3), consider the determinant of $X_{i}$ as computed via Laplace expansion along the row indexed by $j$:
@@ 4. Resolving the Cokernel over $Q$
$0\to\operatorname{Der}_{R\mid k}\to R^{n(n+1)}\overset{J^{T}_{R}}{\to}R^{n+1}\to\operatorname{coker}J^{T}_{R}\to 0$
@@
$Q^{(n-1)(n+1)}\overset{\partial_{2}}{\longrightarrow}Q^{n(n+1)}\overset{J^{T}_{Q}}{\longrightarrow}Q^{n+1}$
@@
$0\longrightarrow F_{s}\overset{d_{s}}{\longrightarrow}F_{s-1}\longrightarrow\ldots\longrightarrow F_{1}\overset{d_{1}}{\longrightarrow}F_{0}$
@@
$0\longrightarrow Q^{(n-1)(n+1)}\overset{\partial_{2}}{\longrightarrow}Q^{n(n+1)}\overset{J^{T}_{Q}}{\longrightarrow}Q^{n+1}\longrightarrow\operatorname{coker}J^{T}_{R}\longrightarrow 0$
@@ 5. Differential Graded Structure
$0\longrightarrow\underset{\{T_{i}\}}{Q^{n}}\overset{\varphi_{2}}{\longrightarrow}\underset{\{e_{i}\}}{Q^{n+1}}\overset{\varphi_{1}}{\longrightarrow}\underset{\{1\}}{Q}\longrightarrow R\longrightarrow 0,$
@@
$0\longrightarrow\underset{\{c_{i,j}\}}{Q^{(n-1)(n+1)}}\overset{\partial_{2}}{\longrightarrow}\underset{\{b_{j,k}\}}{Q^{n(n+1)}}\overset{J^{T}_{Q}}{\longrightarrow}\underset{\{a_{i}\}}{Q^{n+1}}\longrightarrow\operatorname{coker}J^{T}_{R}\longrightarrow 0$
@@
To see that
$\sum_{\ell\neq j}\sum_{r\neq i}\left((-1)^{r}\dfrac{\partial F_{r}}{\partial x_{j,i}}x_{\ell,r}b_{\ell,i}\right)=0,$
we examine
$\sum_{r\neq i}\left((-1)^{r}\dfrac{\partial F_{r}}{\partial x_{j,i}}x_{\ell,r}b_{\ell,i}\right)$
for a fixed $\ell\neq j$. Let $Y$ be the matrix obtained from $X_{i}$ by replacing the row indexed by $j$ with a second row indexed by $\ell$. By Laplace expansion along the row in $Y$ originally indexed by $\ell$, we get
When $i>r$,
$\dfrac{\partial F_{r}}{\partial x_{j,i}}=(-1)^{i+j+1}\det X_{i,r}^{j},$
and when $i<r$,
$\dfrac{\partial F_{r}}{\partial x_{j,i}}=(-1)^{i+j}\det X_{i,r}^{j}.$
So,
Thus,
$\sum_{\ell\neq j}\sum_{r\neq i}\left((-1)^{r}\dfrac{\partial F_{r}}{\partial x_{j,i}}x_{\ell,r}b_{\ell,i}\right)=0.$
@@
$-\sum_{r\neq i}(-1)^{r}\dfrac{\partial F_{r}}{\partial x_{1,k}}x_{\ell,r}+(-1)^{i+1}\dfrac{\partial F_{i}}{\partial x_{1,k}}x_{\ell,i}=-\left(\sum_{r=1}^{n+1}(-1)^{r}\dfrac{\partial F_{r}}{\partial x_{1,k}}x_{\ell,r}\right).$
Let $Y$ be the matrix obtained from $X_{k}$ by replacing the row indexed 1 by a second row indexed $\ell$. The determinant of $Y$ is zero, and one way to compute $\det Y$ is via Laplace expansion along the first row:
$0=\det Y$ $=(-1)^{\ell}\left(\sum_{r<k}(-1)^{r}x_{\ell,r}\det X^{1}_{r,k}+\sum_{r>k}(-1)^{r+1}x_{\ell,r}\det X^{1}_{r,k}\right)$
$=(-1)^{\ell}\left(\sum_{r<k}(-1)^{r+1}x_{\ell,r}\dfrac{\partial F_{r}}{\partial x_{1,k}}+\sum_{r>k}(-1)^{r+1}x_{\ell,r}\dfrac{\partial F_{r}}{\partial x_{1,k}}\right)$
$=(-1)^{\ell+1}\left(\sum_{r<k}(-1)^{r}x_{\ell,r}\dfrac{\partial F_{r}}{\partial x_{1,k}}+\sum_{r>k}(-1)^{r}x_{\ell,r}\dfrac{\partial F_{r}}{\partial x_{1,k}}\right)$
$=\sum_{r=1}^{n+1}(-1)^{r}\dfrac{\partial F_{r}}{\partial x_{1,k}}x_{\ell,r}.$
@@
$\sum_{r=1}^{n+1}\dfrac{\partial F_{r}}{\partial x_{j,k}}(-1)^{r}x_{u,r}=0.$
To see this, let $Y$ be the matrix obtained from $X_{k}$ by replacing the row indexed by $j$ by a second row indexed by $u$. The determinant of $Y$ is zero, and one way to compute $\det Y$ is via Laplace expansion along one of the rows indexed by $u$:
$0=\det Y$ =±$=\pm\left(\sum_{r<k}(-1)^{u+r}x_{u,r}\det X_{k,r}^{u}+\sum_{r>k}(-1)^{u+r+1}x_{u,r}\det X_{k,r}^{u}\right)$
(Remark 3.2) =±$=\pm\left(\sum_{r<k}(-1)^{u+r}x_{u,r}(-1)^{u+k+1}\dfrac{\partial F_{r}}{\partial x_{u,k}}+\sum_{r>k}(-1)^{u+r+1}x_{u,r}(-1)^{u+k}\dfrac{\partial F_{r}}{\partial x_{u,k}}\right)$
=±$=\pm(-1)^{k+1}\sum_{r=1}^{n+1}(-1)^{r}x_{u,r}\dfrac{\partial F_{r}}{\partial x_{j,k}}.$
@@ 6. The Resolution of $\operatorname{Der}_{R\mid k}$
$\dots\longrightarrow R^{n(n+1)(n+2)}\overset{\small\begin{bmatrix}J^{T}_{R}&\varphi_{2}\\ M_{1,1}&M_{2,0}\end{bmatrix}}{\longrightarrow}R^{2n(n+1)}\overset{\small\begin{bmatrix}\partial_{2}&M_{1,0}\end{bmatrix}}{\longrightarrow}\normalsize R^{n(n+1)}\overset{J^{T}_{R}}{\longrightarrow}R^{n+1}\longrightarrow\operatorname{coker}J^{T}_{R}\longrightarrow 0.$
@@
$\dots\longrightarrow R^{n(n+1)(n+2)}\overset{\small\begin{bmatrix}J^{T}_{R}&\varphi_{2}\\ M_{1,1}&M_{2,0}\end{bmatrix}}{\longrightarrow}R^{2n(n+1)}\overset{\small\begin{bmatrix}\partial_{2}&M_{1,0}\end{bmatrix}}{\longrightarrow}\operatorname{Der}_{R\mid k}\longrightarrow 0.$
Theorem 6.3 recovers a minimal generating set for $\operatorname{Der}_{R\mid k}$ (see [bv] for another minimal generating set of $\operatorname{Der}_{R\mid k}$, obtained using a different approach).
@@
To obtain $P^{R}_{\operatorname{Der}_{R\mid k}}(t)$, multiply by $\dfrac{1}{t^{2}}$. ∎
@@
Acknowledgements
Special thanks to Claudia Miller for all of her guidance on this project. Thanks also to Josh Pollitz and Jack Jeffries for helpful discussions. Several results in this paper I first conjectured based on examples I computed using Macaulay2 [M2].

### v2

@@ Abstract.
Key words and phrases:
resolution, derivations, determinantal ring
@@ 1. Introduction
This paper is concerned with showing that a truncation of a relative bar resolution minimally resolves the module of $k$-linear derivations $\operatorname{Der}_{R\mid k}$ on a determinantal ring $R$ defined by the maximal minors of an $n\times(n+1)$ generic matrix $X$. The minimality of the resolution yields a minimal generating set for $\operatorname{Der}_{R\mid k}$ over $R$, a result which may be found in [bv] or [CB], each produced by a starkly different methods.
While minimal generators of modules of derivations on determinantal rings are known, the minimal graded free resolution of $\operatorname{Der}_{R\mid k}$ over $R$ has not been computed. The first order differential operators $D^{1}_{R\mid k}$ on $R$ decompose as $D^{1}_{R\mid k}\cong\operatorname{Der}_{R\mid k}\oplus R$, and so after obtaining the resolution of $\operatorname{Der}_{R\mid k}$, the resolution of $D^{1}_{R\mid k}$ is immediate. Known resolutions of differential operators are few -- see [rc] for the case of low-order differential operators on an isolated hypersurface singularity and [yu] for the case of derivations on a generic hyperplane arrangement, as two examples. Differential operators need not behave the same in positive characteristic as in characteristic zero, so it seems worthwhile to point out that our approach to the minimal graded $R$-free resolution of $\operatorname{Der}_{R\mid k}$ is characteristic-free.
@@
$0\to\operatorname{Der}_{R\mid k}\to R^{n(n+1)}\xrightarrow{J^{T}_{R}}R^{n+1}\to\operatorname{coker}J^{T}_{R}\to 0$
@@
$\dots\longrightarrow R^{n(n+1)(n+2)}\xrightarrow{\small\begin{bmatrix}J^{T}_{R}&\varphi_{2}\\ M_{1,1}&M_{2,0}\end{bmatrix}}R^{2n(n+1)}\xrightarrow{\small\begin{bmatrix}\partial_{2}&M_{1,0}\end{bmatrix}}\operatorname{Der}_{R\mid k}\longrightarrow 0,$
@@ 2. Background
$0\longrightarrow Q^{n}\xrightarrow{\varphi_{2}}Q^{n+1}\xrightarrow{\varphi_{1}}Q\longrightarrow R\longrightarrow 0,$
@@
The ring $R$ is Golod if and only if $\mathcal{A}_{\geq 1}\cdot\mathcal{A}_{\geq 1}\subseteq\mathfrak{m}_{Q}\mathcal{A}$, where $\mathfrak{m}_{Q}$ is the (homogeneous) maximal ideal of $Q$. Similarly, an $R$-module $N$ is Golod if and only if its minimal $Q$-free resolution $\mathcal{U}$ is a dg module over $\mathcal{A}$ with structure satisfying that the action of $\mathcal{A}_{\geq 1}$ on $\mathcal{U}$ lands in $\mathfrak{m}_{Q}\mathcal{U}$.
@@
where $\overline{A}$ is the complex of $S$-modules $\operatorname{coker}(S\to A)$. In our setting, where $A_{0}=S$, this says that $\overline{A}=\Sigma A_{\geq 1}$, the shift of the positive part of $A$ by one homological degree.
@@
$\partial(s\otimes a_{1}\otimes\ldots\otimes a_{p}\otimes u)=$$\pm\sum_{k=1}^{p}s\otimes a_{1}\otimes\ldots\otimes\partial^{A}(a_{k})\ldots\otimes a_{p}\otimes u$
±$\pm s\otimes a_{1}\otimes\ldots\otimes a_{p}\otimes\partial^{U}(u)$
±$\pm\sum_{k=1}^{p-1}s\otimes a_{1}\otimes\ldots\otimes a_{k}\cdot a_{k+1}\otimes\ldots\otimes a_{p}\otimes u$
±$\pm s\otimes a_{1}\otimes\ldots\otimes a_{p}\cdot u,$
where $s\in S/J$, $a_{\ell}\in\overline{A_{i_{\ell}}}$, and $u\in U_{j}$. Exact signs are suppressed, and they are unimportant to the moral of the story.
Remark 2.10.
One can define the relative bar resolution more generally using $A_{\infty}$-structures (see, for example, [Burke]). For our results, we need only the definition in the dg algebra setting.
@@ A $k$-linear derivation $\delta$ on the polynomial $k$-algebra $k[x_{1},\ldots,x_{s}]$ can be written
where $\alpha_{j}\in k[x_{1},\ldots,x_{s}]$ and $\dfrac{\partial}{\partial x_{j}}$ is partial differentiation with respect to $x_{j}$. The prototypical example of a derivation is the Euler derivation
$E:=\sum_{j=1}^{s}x_{j}\dfrac{\partial}{\partial x_{j}},$
which satisfies $E(f):=\deg(f)\cdot f$ for homogeneous $f$.
@@ 3. Useful Lemmas
(3) $\sum_{r\neq i}(-1)^{r}\dfrac{\partial F_{r}}{\partial x_{j,i}}x_{\ell,r}=\begin{cases}0,&j\neq\ell,\\ (-1)^{i+1}F_{i},&j=\ell,\end{cases}$
@@
To see the case of (3) when $j\neq\ell$, let $Y$ be the matrix obtained from $X_{i}$ by replacing the row indexed by $j$ with a second row indexed by $\ell$. By Laplace expansion along the row in $Y$ originally indexed by $\ell$, we get
When $i>r$,
$\dfrac{\partial F_{r}}{\partial x_{j,i}}=(-1)^{i+j+1}\det X_{i,r}^{j},$
and when $i<r$,
$\dfrac{\partial F_{r}}{\partial x_{j,i}}=(-1)^{i+j}\det X_{i,r}^{j}.$
So,
Thus,
$\sum_{r\neq i}\left((-1)^{r}\dfrac{\partial F_{r}}{\partial x_{j,i}}x_{\ell,r}b_{\ell,i}\right)=0.$
To see the case of (3) when $j=\ell$, consider the determinant of $X_{i}$ as computed via Laplace expansion along the row indexed by $j$:
@@ 4. Resolving the Cokernel over $Q$
$0\to\operatorname{Der}_{R\mid k}\to R^{n(n+1)}\xrightarrow{J^{T}_{R}}R^{n+1}\to\operatorname{coker}J^{T}_{R}\to 0$
@@
$Q^{(n-1)(n+1)}\xrightarrow{\partial_{2}}Q^{n(n+1)}\xrightarrow{J^{T}_{Q}}Q^{n+1}$
@@
$0\longrightarrow F_{s}\xrightarrow{d_{s}}F_{s-1}\longrightarrow\ldots\longrightarrow F_{1}\xrightarrow{d_{1}}F_{0}$
@@
$0\longrightarrow Q^{(n-1)(n+1)}\xrightarrow{\partial_{2}}Q^{n(n+1)}\xrightarrow{J^{T}_{Q}}Q^{n+1}\longrightarrow\operatorname{coker}J^{T}_{R}\longrightarrow 0$
@@ 5. Differential Graded Structure
$0\longrightarrow\underset{\{T_{i}\}}{Q^{n}}\xrightarrow{\varphi_{2}}\underset{\{e_{i}\}}{Q^{n+1}}\xrightarrow{\varphi_{1}}\underset{\{1\}}{Q}\longrightarrow R\longrightarrow 0,$
@@
$0\longrightarrow\underset{\{c_{i,j}\}}{Q^{(n-1)(n+1)}}\xrightarrow{\partial_{2}}\underset{\{b_{j,k}\}}{Q^{n(n+1)}}\xrightarrow{J^{T}_{Q}}\underset{\{a_{i}\}}{Q^{n+1}}\longrightarrow\operatorname{coker}J^{T}_{R}\longrightarrow 0$
@@
$-\sum_{r\neq i}(-1)^{r}\dfrac{\partial F_{r}}{\partial x_{1,k}}x_{\ell,r}+(-1)^{i+1}\dfrac{\partial F_{i}}{\partial x_{1,k}}x_{\ell,i}=-\left(\sum_{r=1}^{n+1}(-1)^{r}\dfrac{\partial F_{r}}{\partial x_{1,k}}x_{\ell,r}\right)=0,$
by Lemma 3.8, equation (3), and Remark 2.15.
@@
$\sum_{r=1}^{n+1}\dfrac{\partial F_{r}}{\partial x_{j,k}}(-1)^{r}x_{u,r}=0,$
by Lemma 3.8, equation (3), and Remark 2.15.
@@ 6. The Resolution of $\operatorname{Der}_{R\mid k}$
$\dots\longrightarrow R^{n(n+1)(n+2)}\xrightarrow{\small\begin{bmatrix}J^{T}_{R}&\varphi_{2}\\ M_{1,1}&M_{2,0}\end{bmatrix}}R^{2n(n+1)}\xrightarrow{\small\begin{bmatrix}\partial_{2}&M_{1,0}\end{bmatrix}}\normalsize R^{n(n+1)}\xrightarrow{J^{T}_{R}}R^{n+1}\longrightarrow\operatorname{coker}J^{T}_{R}\longrightarrow 0.$
@@
$\dots\longrightarrow R^{n(n+1)(n+2)}\xrightarrow{\small\begin{bmatrix}J^{T}_{R}&\varphi_{2}\\ M_{1,1}&M_{2,0}\end{bmatrix}}R^{2n(n+1)}\xrightarrow{\small\begin{bmatrix}\partial_{2}&M_{1,0}\end{bmatrix}}\operatorname{Der}_{R\mid k}\longrightarrow 0.$
Theorem 6.3 recovers a minimal generating set for $\operatorname{Der}_{R\mid k}$ (see [bv] and [CB] for other minimal generating sets of $\operatorname{Der}_{R\mid k}$, obtained using different approaches).
@@
By the exact sequence
$0\to\operatorname{Der}_{R\mid k}\to R^{n(n+1)}\xrightarrow{J^{T}_{R}}R^{n+1}\to\operatorname{coker}J^{T}_{R}\to 0,$
we have that $P^{R}_{\operatorname{Der}_{R\mid k}}(t)$ is obtained by subtracting $(n+1)+n(n+1)t$ and multiplying by $\dfrac{1}{t^{2}}$:
$P^{R}_{\operatorname{Der}_{R\mid k}}(t)=\left(\dfrac{2nt^{2}+2n^{2}t^{2}+n^{2}t^{3}+n^{3}t^{3}}{1-t-nt^{2}}\right)\cdot\dfrac{1}{t^{2}}=\dfrac{n(n+1)(2+nt)}{1-t-nt^{2}}.$
@@
Acknowledgments
Special thanks to Claudia Miller for all of her guidance on this project. Thanks also to Josh Pollitz and Jack Jeffries for helpful discussions and to the reviewers for their comments and suggestions. Several results in this paper I first conjectured based on examples I computed using Macaulay2 [M2].